# SVFAP: Self-supervised Video Facial Affect Perceiver

### v1

@@ 4.2.4 Comparisons with State-of-the-art Methods
We first show the comparison results on three large in-the-wild datasets, including DFEW, FERV39k, and MAFW. The results on three datasets are reported in Table VII, VIII, and IX, respectively. The comparison baselines can be roughly divided into three categories: 1) the combination of convolution neural network (CNN) and recurrent neural network (RNN), i.e., CNN+RNN, such as ResNet+LSTM [30, 31]. 2) classic 3D CNNs, including 3D ResNet [38], C3D [35], P3D [37], and R(2+1)D [72]. 3) hybrid architectures of CNN and Transformer, e.g., Former-DFER [12], STT [13], NR-DFERNet [71], GCA+IAL, and T-ESFL [34].
As shown in Table VII, we observe that SVFAP-B outperforms previous state-of-the-art methods on DFEW significantly (i.e., 5.72% UAR and 5.03% WAR improvement), setting a new record on this dataset. Besides, the small version, SVFAP-S, also surpasses the best-performing methods by a large margin, achieving better accuracy-complexity trade-off. When comparing the fine-grained performance of each class, we find that our methods achieve remarkable improvements for most emotions (e.g., happy and sad). Notably, for the rare disgust emotion which only accounts for 1.2% (about 146 samples) in the whole dataset, most baseline methods fail to classify its samples correctly. Nevertheless, our SVFAP-B improves the previous best performer by about 10%. This result demonstrates that the proposed method can learn generic affect-related representations via large-scale self-supervised pre-training, thus alleviating the unbalanced learning in minority classes. Moreover, we have similar observations on the other two datasets. On the largest DFER dataset FERV39k, as shown in Table VIII, SVFAP-B achieves 42.14% UAR and 52.29% WAR, outperforming the best baselines by 4.38% UAR and 3.85% WAR. On the MAFW dataset, as given in Table IX, SVFAP-B improves over the-state-of-the-art T-ESFL by 7.91% UAR and 6.10% WAR. Besides, a slight performance drop is also observed for both datasets. To sum up, the above encouraging results on three in-the-wild datasets verify the strong generalization ability of our SVFAP in real-world scenarios.
Finally, we present the comparison results on three small lab-controlled datasets, including CREMA-D, RAVDESS, and eNTERFACE05. The results are reported in Table X, XI, and XII, respectively. Similarly, we observe consistently significant performance improvements on these datasets. For instance, as shown in Table X, SVFAP-B outperforms the best-performing unimodal methods by about 12% UAR and 10% WAR on CREMA-D. We also report the results of several multimodal methods. Compared with them, SVFAP-B still show slightly better performance without using the audio information, which demonstrates the overwhelming superiority of the proposed method again.
TABLE X: Comparison with state-of-the-art methods on CREMA-D.
@@
TABLE XI: Comparison with state-of-the-art methods on RAVDESS.
@@
TABLE XII: Comparison with state-of-the-art methods on eNTERFACE05.
@@ 4.3 Dimensional Emotion Recognition
TABLE XIII: Comparison with state-of-the-art methods on Werewolf-XL.
@@
TABLE XIV: Comparison with state-of-the-art methods on AVCAffe.
@@
4.3.1 Datasets
Werewolf-XL [87] is a spontaneous audio-visual dataset with a total of 890 minutes of videos recorded during competitive group interactions in Werewolf games. It contains 131,688 video clips from 129 subjects in 30 game sessions, including 14,632 samples from active speakers and the rest for listeners in the game. Werewolf-XL provides both self-reported categorical emotion labels and external assessed dimensional emotion scores. In this paper, we only use 14,632 speaker samples and the dimensional annotations. As the goal is to predict continuous scores, we formulate it as a regression problem as stated in Section 3.2. Besides, we adopt a 5-fold session-independent cross-validation protocol for evaluation. Two type of standard metrics, i.e., Concordance Correlation Coefficient (CCC) and Pearson Correlation Coefficient (PCC), are reported in this paper. They are computed as follows:
AVCAffe [88] is currently the largest audio-visual dataset with both affect and cognitive load attributes. It is recorded by simulating remote work setting and contains more than 108 hours of video from 106 subjects. Each subject is asked to participate 7 specially designed tasks. After each task, self-reported affect (i.e., arousal and valence) and cognitive load scores are collected as ground truth labels. We only predict arousal and valence scores on a scale of 0-4 and formulate it as a classification problem according to the original paper. Since the duration of each task are too long (7.5 minutes), each video has been segmented into multiple 6-second short clips for model training. We follow the paper to obtain video-level predictions by averaging clip-level scores and employ the weighted F1-score as the evaluation metric. The dataset provides official split: 86 subjects for training and 20 subjects for test.
@@ 4.3.2 Comparison with State-of-the-art Methods
The results on Werewolf-XL are shown in Table XIII. As we can see, our method achieves much better performance than unimodal baseline methods in three emotion dimensions. Specifically, SVFAP-B outperforms the best-performing ones by about 0.04 CCC and 0.03 PCC in arousal, 0.04 CCC and 0.04 PCC in valence, and 0.13 CCC and 0.10 PCC in dominance. Besides, we also show the multimodal baseline in the table. Compared with it, SVFAP-B still shows superior performance in valence, although worse results are achieved in arousal and dominance. Besides, we only observe a moderate performance degradation for SVFAP-S on this dataset.
We further present the comparison results on AVCAffe in Table XIV. Similarly, we observe that both SVFAP-B and SVFAP-S surpass the state-of-the-art unimodal methods by a large margin (especially in arousal). For instance, SVFAP-B achieves about 6% F1-score improvement in arousal and about 2% in valence. Moreover, when compared with multimodal methods, our base model still presents a comparable performance in both arousal and valence, which verifies the effectiveness of large-scale self-supervised pre-training for dimension emotion recognition.
@@ 4.4 Personality Recognition
TABLE XV: Comparison with state-of-the-art methods on ChaLearn 2016 First Impression in terms of CCC. O: Openness. C: Conscientiousness. E: Extraversion. A: Agreeableness. N: Neuroticism.
@@
TABLE XVI: Comparison with state-of-the-art methods on ChaLearn 2016 First Impression in terms of ACC. O: Openness. C: Conscientiousness. E: Extraversion. A: Agreeableness. N: Neuroticism.
@@
The classic ChaLearn First Impression dataset [101] is selected for evaluation. It consists of 10,000 talking-to-the-camera clips extracted from over 3,000 high-definition YouTube videos. Each video clip has a duration of about 15 seconds and is annotated with the Big-Five personality traits (i.e., openness, conscientiousness, extraversion, agreeableness, and neuroticism). This dataset has been split into three sets: 6000 videos for training, 2000 videos for validation, and 2000 videos for the final test. Since the task is to predict continuous scores in different personality dimensions, we formulate it as a regression problem. Following [102], we use two standard evaluation metrics, i.e., CCC and the Accuracy (ACC). CCC is defined in Eq. (14). The definition of ACC is given as follows:
@@
The comparison results in terms of CCC and ACC are shown in Table XV and Table XVI, respectively. For CCC, we observe that two versions of our models achieve similar performance. And both of them outperform the best-performing supervised baseline largely, achieving about 0.02-0.03 CCC improvement. For ACC, we find that the variation between different methods is relatively small. Nevertheless, our methods still show the best performance.
@@ 5 Conclusion
In this paper, we have presented a self-supervised learning method, termed Self-supervised Video Facial Affect Perceiver (SVFAP), to unleash the power of large-scale self-supervised pre-training for video-based facial affect analysis. SVFAP utilizes masked facial video autoencoding as the objective to perform self-supervised pre-training on a large amount of unlabeled facial videos. Besides, it employs a novel TBSBT model as the encoder to minimize large redundancy in 3D facial video data from both spatial and temporal perspectives, leading to significantly lower computational cost and superior performance. To verify the effectiveness of SVFAP, we conduct extensive experiments on nine datasets in three popular downstream tasks, including dynamic facial expression recognition, dimensional emotion recognition, and personality recognition. The results demonstrate that SVFAP can learn powerful affect-related representations via large-scale self-supervised pre-training. Specifically, it largely outperforms previous pre-trained supervised and self-supervised models and also shows strong adaptation ability in the low data regime. Moreover, our SVFAP achieves significant improvements over state-of-the-art methods in three downstream tasks, setting new records on all datasets.
In future work, we plan to investigate the scaling behavior of SVFAP using larger models and more unlabeled data. Besides, we are interested to evaluate SVFAP on other downstream tasks, such as dynamic micro-expression recognition and depression level detection. We also hope our work can inspire more relevant research to further advance the development of video-based facial affect analysis.
@@ Acknowledgments
This work is supported by the National Natural Science Foundation of China (NSFC) (No.62201572, No.61831022, No.62276259, No.U21B2010), Beijing Municipal Science & Technology Commission, Administrative Commission of Zhongguancun Science Park (No.Z211100004821013), Open Research Projects of Zhejiang Lab (No.2021KH0AB06), CCF-Baidu Open Fund (No.OF2022025).

### v2

@@ 4.2.4 Comparisons with State-of-the-art Methods
We first show the comparison results on three large in-the-wild datasets, including DFEW, FERV39k, and MAFW. The results on three datasets are reported in Table VIII, IX, and X, respectively. The comparison baselines can be roughly divided into three categories: 1) the combination of convolution neural network (CNN) and recurrent neural network (RNN), i.e., CNN+RNN, such as ResNet+LSTM [31, 32]. 2) classic 3D CNNs, including 3D ResNet [39], C3D [36], P3D [38], and R(2+1)D [76]. 3) hybrid architectures of CNN and Transformer, e.g., Former-DFER [13], STT [14], NR-DFERNet [74], IAL [35], and T-ESFL [28].
As shown in Table VIII, we observe that SVFAP-B outperforms previous state-of-the-art methods on DFEW significantly (i.e., 5.72% UAR and 5.02% WAR improvement), setting a new record on this dataset. Besides, the small version, SVFAP-S, also surpasses the best-performing methods by a large margin, achieving a better accuracy-complexity trade-off. When comparing the fine-grained performance of each class, we find that our methods achieve remarkable improvements for most emotions (e.g., happy and sad). Notably, for the rare disgust emotion which only accounts for 1.2% (about 146 samples) in the whole dataset, most baseline methods fail to classify its samples correctly. Nevertheless, our SVFAP-B improves the previous best performer by about 10%. This result demonstrates that the proposed method can learn generic affect-related representations via large-scale self-supervised pre-training, thus alleviating the unbalanced learning in minority classes. Moreover, we have similar observations on the other two datasets. On the largest DFER dataset FERV39k, as shown in Table IX, SVFAP-B achieves 42.14% UAR and 52.29% WAR, outperforming the best baselines by 4.38% UAR and 3.75% WAR. On the MAFW dataset, as given in Table X, SVFAP-B improves over the state-of-the-art T-ESFL by 7.91% UAR and 6.10% WAR. Besides, a slight performance drop is also observed for both datasets. To sum up, the above encouraging results on three in-the-wild datasets verify the strong generalization ability of our SVFAP in real-world scenarios.
Finally, we present the comparison results on three small lab-controlled datasets, including CREMA-D, RAVDESS, and eNTERFACE05. The results are reported in Table XI, XII, and XIII, respectively. Similarly, we observe consistently significant performance improvements on these datasets. For instance, as shown in Table XI, SVFAP-B outperforms the best-performing unimodal methods by about 12% UAR and 10% WAR on CREMA-D. We also report the results of several multimodal methods. Compared with them, SVFAP-B still shows slightly better performance without using the audio information, which demonstrates the overwhelming superiority of the proposed method again.
TABLE XI: Comparison with state-of-the-art methods on CREMA-D.
@@
TABLE XII: Comparison with state-of-the-art methods on RAVDESS.
@@
TABLE XIII: Comparison with state-of-the-art methods on eNTERFACE05.
@@ 4.3 Dimensional Emotion Recognition
4.3.1 Datasets
Werewolf-XL [90] is a spontaneous audio-visual dataset with a total of 890 minutes of videos recorded during competitive group interactions in Werewolf games. It contains 131,688 video clips from 129 subjects in 30 game sessions, including 14,632 samples from active speakers and the rest from listeners in the game. Werewolf-XL provides both self-reported categorical emotion labels and externally assessed dimensional emotion scores. In this paper, we only use 14,632 speaker samples and dimensional annotations. As the goal is to predict continuous scores, we formulate it as a regression problem as stated in Section 3.2. Besides, we adopt a 5-fold session-independent cross-validation protocol for evaluation. Two types of standard metrics, i.e., Concordance Correlation Coefficient (CCC) and Pearson Correlation Coefficient (PCC), are reported in this paper. They are computed as follows:
AVCAffe [91] is currently the largest audio-visual dataset with both affect and cognitive load attributes. It is recorded by simulating a remote work setting and contains more than 108 hours of video from 106 subjects. Each subject is asked to participate in 7 specially designed tasks. After each task, self-reported affect (i.e., arousal and valence) and cognitive load scores are collected as ground truth labels. We only predict arousal and valence scores on a scale of 0-4 and formulate it as a classification problem according to the original paper. Since the duration of each task is too long (7.5 minutes), each video has been segmented into multiple 6-second short clips for model training. We follow the paper to obtain video-level predictions by averaging clip-level scores and employ the weighted F1 score as the evaluation metric. The dataset provides an official split: 86 subjects for training and 20 subjects for test.
TABLE XIV: Comparison with state-of-the-art methods on Werewolf-XL.
@@ 4.3.1 Datasets
TABLE XV: Comparison with state-of-the-art methods on AVCAffe.
@@ 4.3.2 Comparison with State-of-the-art Methods
The results on Werewolf-XL are shown in Table XIV. As we can see, our method achieves much better performance than unimodal baseline methods in three emotion dimensions. Specifically, SVFAP-B outperforms the best-performing ones by about 0.04 CCC and 0.03 PCC in arousal, 0.04 CCC and 0.04 PCC in valence, and 0.13 CCC and 0.10 PCC in dominance. Besides, we also show the multimodal baseline in the table. Compared with it, SVFAP-B still shows superior performance in valence, although worse results are achieved in arousal and dominance. Besides, we only observe a moderate performance degradation for SVFAP-S on this dataset.
We further present the comparison results on AVCAffe in Table XV. Similarly, we observe that both SVFAP-B and SVFAP-S surpass the state-of-the-art unimodal methods by a large margin (especially in arousal). For instance, SVFAP-B achieves about 6% F1-score improvement in arousal and about 2% in valence. Moreover, when compared with multimodal methods, our base model still presents a comparable performance in both arousal and valence, which verifies the effectiveness of large-scale self-supervised pre-training for dimension emotion recognition.
@@ 4.4 Personality Recognition
TABLE XVI: Comparison with state-of-the-art methods on ChaLearn 2016 First Impression in terms of CCC. O: Openness. C: Conscientiousness. E: Extraversion. A: Agreeableness. N: Neuroticism.
@@
TABLE XVII: Comparison with state-of-the-art methods on ChaLearn 2016 First Impression in terms of ACC. O: Openness. C: Conscientiousness. E: Extraversion. A: Agreeableness. N: Neuroticism.
@@
The classic ChaLearn First Impression dataset [105] is selected for evaluation. It consists of 10,000 talking-to-the-camera clips extracted from over 3,000 high-definition YouTube videos. Each video clip has a duration of about 15 seconds and is annotated with the Big Five personality traits (i.e., openness, conscientiousness, extraversion, agreeableness, and neuroticism). This dataset has been split into three sets: 6000 videos for training, 2000 videos for validation, and 2000 videos for the final test. Since the task is to predict continuous scores in different personality dimensions, we formulate it as a regression problem. Following [106], we use two standard evaluation metrics, i.e., CCC and the Accuracy (ACC). CCC is defined in Eq. (14). The definition of ACC is given as follows:
@@
The comparison results in terms of CCC and ACC are shown in Table XVI and Table XVII, respectively. For CCC, we observe that two versions of our models achieve similar performance. Both of them outperform the best-performing supervised baseline largely, achieving about 0.02-0.03 CCC improvement. For ACC, we find that the variation between different methods is relatively small. Nevertheless, our methods still show the best performance.
@@ 5 Conclusion
In this paper, we have presented a self-supervised learning method, termed Self-supervised Video Facial Affect Perceiver (SVFAP), to unleash the power of large-scale self-supervised pre-training for video-based facial affect analysis. SVFAP utilizes masked facial video autoencoding as the objective to perform self-supervised pre-training on a large amount of unlabeled facial videos. Besides, it employs a novel TBSBT model as the encoder to minimize large redundancy in 3D facial video data from both spatial and temporal perspectives, leading to significantly lower computational costs and superior performance. To verify the effectiveness of SVFAP, we conduct extensive experiments on nine datasets in three popular downstream tasks, including dynamic facial expression recognition, dimensional emotion recognition, and personality recognition. The results demonstrate that SVFAP can learn powerful affect-related representations via large-scale self-supervised pre-training. Specifically, it largely outperforms previous pre-trained supervised and self-supervised models and also shows strong adaptation ability in the low data regime. Moreover, our SVFAP achieves significant improvements over state-of-the-art methods in three downstream tasks, setting new records on all datasets.
In future work, we plan to investigate the scaling behavior of SVFAP using larger models and more unlabeled data. Besides, it is also interesting to evaluate SVFAP in other downstream tasks, such as dynamic micro-expression recognition and depression level detection. We also hope our work can inspire more relevant research to further advance the development of video-based facial affect analysis.
@@ Acknowledgments
This work is supported by the National Natural Science Foundation of China (NSFC) ( No.62276259, No.62201572, No.U21B2010, No.62271083, No.62306316).